A group of four items has a mean of 69. What would the value of the fifth item have to be in order to produce a mean of 75?
Answer by TimothyLamb(4379)

(69*4 + x)/5 = 75
(69*4) + x = 75*5
x = 75*5 - 69*4
x = 99
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---